# Credal Prediction based on Relative Likelihood

### v1

@@ C.4 Performance of Ensemble Members
ChaosNLI CreWra CreEnsα CreBNN CreNet Member $\bm{m}$ 1 $0.658\scriptstyle{\pm 0.013}$ $0.671\scriptstyle{\pm 0.011}$ $0.512\scriptstyle{\pm 0.083}$ $0.525\scriptstyle{\pm 0.051}$ 2 $0.648\scriptstyle{\pm 0.012}$ $0.676\scriptstyle{\pm 0.013}$ $0.453\scriptstyle{\pm 0.010}$ $0.488\scriptstyle{\pm 0.098}$ 3 $0.641\scriptstyle{\pm 0.020}$ $0.676\scriptstyle{\pm 0.013}$ $0.506\scriptstyle{\pm 0.085}$ $0.542\scriptstyle{\pm 0.030}$ 4 $0.636\scriptstyle{\pm 0.012}$ $0.666\scriptstyle{\pm 0.008}$ $0.511\scriptstyle{\pm 0.092}$ $0.485\scriptstyle{\pm 0.037}$ 5 $0.651\scriptstyle{\pm 0.032}$ $0.667\scriptstyle{\pm 0.011}$ $0.532\scriptstyle{\pm 0.111}$ $0.554\scriptstyle{\pm 0.051}$ 6 $0.660\scriptstyle{\pm 0.016}$ $0.674\scriptstyle{\pm 0.021}$ $0.503\scriptstyle{\pm 0.081}$ $0.470\scriptstyle{\pm 0.115}$ 7 $0.637\scriptstyle{\pm 0.014}$ $0.677\scriptstyle{\pm 0.009}$ $0.453\scriptstyle{\pm 0.010}$ $0.518\scriptstyle{\pm 0.036}$ 8 $0.667\scriptstyle{\pm 0.016}$ $0.670\scriptstyle{\pm 0.010}$ $0.453\scriptstyle{\pm 0.010}$ $0.446\scriptstyle{\pm 0.030}$ 9 $0.584\scriptstyle{\pm 0.086}$ $0.663\scriptstyle{\pm 0.005}$ $0.453\scriptstyle{\pm 0.010}$ $0.462\scriptstyle{\pm 0.113}$ 10 $0.663\scriptstyle{\pm 0.035}$ $0.674\scriptstyle{\pm 0.029}$ $0.664\scriptstyle{\pm 0.012}$ $0.531\scriptstyle{\pm 0.064}$ 11 $0.648\scriptstyle{\pm 0.017}$ $0.653\scriptstyle{\pm 0.015}$ $0.591\scriptstyle{\pm 0.099}$ $0.484\scriptstyle{\pm 0.047}$ 12 $0.660\scriptstyle{\pm 0.016}$ $0.662\scriptstyle{\pm 0.016}$ $0.587\scriptstyle{\pm 0.097}$ $0.549\scriptstyle{\pm 0.071}$ 13 $0.664\scriptstyle{\pm 0.014}$ $0.655\scriptstyle{\pm 0.017}$ $0.581\scriptstyle{\pm 0.099}$ $0.487\scriptstyle{\pm 0.040}$ 14 $0.650\scriptstyle{\pm 0.011}$ $0.654\scriptstyle{\pm 0.006}$ $0.655\scriptstyle{\pm 0.007}$ $0.505\scriptstyle{\pm 0.039}$ 15 $0.635\scriptstyle{\pm 0.038}$ $0.651\scriptstyle{\pm 0.018}$ $0.602\scriptstyle{\pm 0.097}$ $0.458\scriptstyle{\pm 0.143}$ 16 $0.652\scriptstyle{\pm 0.026}$ $0.654\scriptstyle{\pm 0.007}$ $0.629\scriptstyle{\pm 0.046}$ $0.486\scriptstyle{\pm 0.012}$ 17 $0.628\scriptstyle{\pm 0.019}$ $0.633\scriptstyle{\pm 0.015}$ $0.587\scriptstyle{\pm 0.086}$ $0.518\scriptstyle{\pm 0.057}$ 18 $0.652\scriptstyle{\pm 0.009}$ $0.605\scriptstyle{\pm 0.021}$ $0.502\scriptstyle{\pm 0.079}$ $0.519\scriptstyle{\pm 0.031}$ 19 $0.660\scriptstyle{\pm 0.012}$ $0.581\scriptstyle{\pm 0.037}$ $0.584\scriptstyle{\pm 0.095}$ $0.487\scriptstyle{\pm 0.050}$ 20 $0.651\scriptstyle{\pm 0.013}$ $0.480\scriptstyle{\pm 0.074}$ $0.586\scriptstyle{\pm 0.095}$ $0.560\scriptstyle{\pm 0.043}$ CIFAR-10 CreWra CreEnsα CreBNN CreNet Member $\bm{m}$ 1 $0.942\scriptstyle{\pm 0.001}$ $0.955\scriptstyle{\pm 0.000}$ $0.875\scriptstyle{\pm 0.004}$ $0.941\scriptstyle{\pm 0.000}$ 2 $0.944\scriptstyle{\pm 0.001}$ $0.954\scriptstyle{\pm 0.000}$ $0.877\scriptstyle{\pm 0.005}$ $0.943\scriptstyle{\pm 0.000}$ 3 $0.944\scriptstyle{\pm 0.001}$ $0.952\scriptstyle{\pm 0.001}$ $0.867\scriptstyle{\pm 0.015}$ $0.944\scriptstyle{\pm 0.001}$ 4 $0.943\scriptstyle{\pm 0.002}$ $0.952\scriptstyle{\pm 0.000}$ $0.881\scriptstyle{\pm 0.003}$ $0.942\scriptstyle{\pm 0.001}$ 5 $0.943\scriptstyle{\pm 0.001}$ $0.952\scriptstyle{\pm 0.001}$ $0.872\scriptstyle{\pm 0.004}$ $0.942\scriptstyle{\pm 0.001}$ 6 $0.943\scriptstyle{\pm 0.000}$ $0.952\scriptstyle{\pm 0.000}$ $0.869\scriptstyle{\pm 0.006}$ $0.942\scriptstyle{\pm 0.002}$ 7 $0.941\scriptstyle{\pm 0.001}$ $0.952\scriptstyle{\pm 0.001}$ $0.877\scriptstyle{\pm 0.001}$ $0.943\scriptstyle{\pm 0.001}$ 8 $0.942\scriptstyle{\pm 0.001}$ $0.951\scriptstyle{\pm 0.001}$ $0.880\scriptstyle{\pm 0.006}$ $0.943\scriptstyle{\pm 0.002}$ 9 $0.943\scriptstyle{\pm 0.001}$ $0.951\scriptstyle{\pm 0.000}$ $0.879\scriptstyle{\pm 0.005}$ $0.944\scriptstyle{\pm 0.001}$ 10 $0.943\scriptstyle{\pm 0.000}$ $0.953\scriptstyle{\pm 0.000}$ $0.873\scriptstyle{\pm 0.007}$ $0.943\scriptstyle{\pm 0.003}$ 11 $0.942\scriptstyle{\pm 0.001}$ $0.953\scriptstyle{\pm 0.001}$ $0.882\scriptstyle{\pm 0.004}$ $0.943\scriptstyle{\pm 0.000}$ 12 $0.943\scriptstyle{\pm 0.000}$ $0.953\scriptstyle{\pm 0.000}$ $0.879\scriptstyle{\pm 0.003}$ $0.942\scriptstyle{\pm 0.000}$ 13 $0.943\scriptstyle{\pm 0.000}$ $0.953\scriptstyle{\pm 0.001}$ $0.874\scriptstyle{\pm 0.000}$ $0.943\scriptstyle{\pm 0.002}$ 14 $0.945\scriptstyle{\pm 0.000}$ $0.952\scriptstyle{\pm 0.001}$ $0.872\scriptstyle{\pm 0.006}$ $0.941\scriptstyle{\pm 0.001}$ 15 $0.942\scriptstyle{\pm 0.002}$ $0.948\scriptstyle{\pm 0.000}$ $0.856\scriptstyle{\pm 0.019}$ $0.942\scriptstyle{\pm 0.001}$ 16 $0.942\scriptstyle{\pm 0.000}$ $0.941\scriptstyle{\pm 0.000}$ $0.870\scriptstyle{\pm 0.002}$ $0.942\scriptstyle{\pm 0.001}$ 17 $0.942\scriptstyle{\pm 0.001}$ $0.930\scriptstyle{\pm 0.001}$ $0.854\scriptstyle{\pm 0.035}$ $0.942\scriptstyle{\pm 0.002}$ 18 $0.944\scriptstyle{\pm 0.001}$ $0.921\scriptstyle{\pm 0.001}$ $0.872\scriptstyle{\pm 0.007}$ $0.943\scriptstyle{\pm 0.002}$ 19 $0.943\scriptstyle{\pm 0.001}$ $0.906\scriptstyle{\pm 0.002}$ $0.873\scriptstyle{\pm 0.001}$ $0.942\scriptstyle{\pm 0.001}$ 20 $0.942\scriptstyle{\pm 0.001}$ $0.872\scriptstyle{\pm 0.002}$ $0.860\scriptstyle{\pm 0.012}$ $0.942\scriptstyle{\pm 0.001}$ QualityMRI CreWra CreEnsα CreBNN CreNet Member $\bm{m}$ 1 $0.457\scriptstyle{\pm 0.099}$ $0.430\scriptstyle{\pm 0.170}$ $0.548\scriptstyle{\pm 0.115}$ $0.581\scriptstyle{\pm 0.137}$ 2 $0.452\scriptstyle{\pm 0.139}$ $0.409\scriptstyle{\pm 0.163}$ $0.484\scriptstyle{\pm 0.115}$ $0.559\scriptstyle{\pm 0.112}$ 3 $0.425\scriptstyle{\pm 0.124}$ $0.419\scriptstyle{\pm 0.160}$ $0.468\scriptstyle{\pm 0.126}$ $0.570\scriptstyle{\pm 0.110}$ 4 $0.419\scriptstyle{\pm 0.126}$ $0.446\scriptstyle{\pm 0.187}$ $0.548\scriptstyle{\pm 0.115}$ $0.575\scriptstyle{\pm 0.102}$ 5 $0.398\scriptstyle{\pm 0.135}$ $0.387\scriptstyle{\pm 0.139}$ $0.468\scriptstyle{\pm 0.126}$ $0.548\scriptstyle{\pm 0.126}$ 6 $0.403\scriptstyle{\pm 0.117}$ $0.398\scriptstyle{\pm 0.158}$ $0.554\scriptstyle{\pm 0.118}$ $0.554\scriptstyle{\pm 0.084}$ 7 $0.409\scriptstyle{\pm 0.175}$ $0.414\scriptstyle{\pm 0.161}$ $0.446\scriptstyle{\pm 0.107}$ $0.554\scriptstyle{\pm 0.095}$ 8 $0.430\scriptstyle{\pm 0.119}$ $0.441\scriptstyle{\pm 0.182}$ $0.554\scriptstyle{\pm 0.118}$ $0.559\scriptstyle{\pm 0.133}$ 9 $0.392\scriptstyle{\pm 0.137}$ $0.419\scriptstyle{\pm 0.181}$ $0.473\scriptstyle{\pm 0.110}$ $0.586\scriptstyle{\pm 0.141}$ 10 $0.446\scriptstyle{\pm 0.177}$ $0.430\scriptstyle{\pm 0.140}$ $0.462\scriptstyle{\pm 0.130}$ $0.570\scriptstyle{\pm 0.118}$ 11 $0.425\scriptstyle{\pm 0.137}$ $0.414\scriptstyle{\pm 0.153}$ $0.554\scriptstyle{\pm 0.118}$ $0.602\scriptstyle{\pm 0.107}$ 12 $0.398\scriptstyle{\pm 0.122}$ $0.430\scriptstyle{\pm 0.153}$ $0.554\scriptstyle{\pm 0.118}$ $0.565\scriptstyle{\pm 0.103}$ 13 $0.398\scriptstyle{\pm 0.130}$ $0.430\scriptstyle{\pm 0.140}$ $0.446\scriptstyle{\pm 0.107}$ $0.575\scriptstyle{\pm 0.090}$ 14 $0.409\scriptstyle{\pm 0.066}$ $0.430\scriptstyle{\pm 0.142}$ $0.457\scriptstyle{\pm 0.122}$ $0.581\scriptstyle{\pm 0.070}$ 15 $0.409\scriptstyle{\pm 0.146}$ $0.414\scriptstyle{\pm 0.132}$ $0.468\scriptstyle{\pm 0.126}$ $0.624\scriptstyle{\pm 0.133}$ 16 $0.414\scriptstyle{\pm 0.107}$ $0.425\scriptstyle{\pm 0.167}$ $0.554\scriptstyle{\pm 0.118}$ $0.608\scriptstyle{\pm 0.100}$ 17 $0.419\scriptstyle{\pm 0.139}$ $0.414\scriptstyle{\pm 0.119}$ $0.554\scriptstyle{\pm 0.118}$ $0.581\scriptstyle{\pm 0.091}$ 18 $0.403\scriptstyle{\pm 0.168}$ $0.414\scriptstyle{\pm 0.129}$ $0.468\scriptstyle{\pm 0.126}$ $0.581\scriptstyle{\pm 0.117}$ 19 $0.425\scriptstyle{\pm 0.151}$ $0.414\scriptstyle{\pm 0.073}$ $0.554\scriptstyle{\pm 0.118}$ $0.565\scriptstyle{\pm 0.142}$ 20 $0.430\scriptstyle{\pm 0.154}$ $0.446\scriptstyle{\pm 0.068}$ $0.548\scriptstyle{\pm 0.117}$ $0.581\scriptstyle{\pm 0.103}$

### v2

@@ C.4 Performance of Ensemble Members
ChaosNLI CreWra CreEnsα CreBNN CreNet Member $\bm{m}$ 1 $0.658\scriptstyle{\pm 0.013}$ $0.671\scriptstyle{\pm 0.011}$ $0.680\scriptstyle{\pm 0.013}$ $0.525\scriptstyle{\pm 0.051}$ 2 $0.648\scriptstyle{\pm 0.012}$ $0.676\scriptstyle{\pm 0.013}$ $0.674\scriptstyle{\pm 0.031}$ $0.488\scriptstyle{\pm 0.098}$ 3 $0.641\scriptstyle{\pm 0.020}$ $0.676\scriptstyle{\pm 0.013}$ $0.517\scriptstyle{\pm 0.079}$ $0.542\scriptstyle{\pm 0.030}$ 4 $0.636\scriptstyle{\pm 0.012}$ $0.666\scriptstyle{\pm 0.008}$ $0.683\scriptstyle{\pm 0.030}$ $0.485\scriptstyle{\pm 0.037}$ 5 $0.651\scriptstyle{\pm 0.032}$ $0.667\scriptstyle{\pm 0.011}$ $0.533\scriptstyle{\pm 0.112}$ $0.554\scriptstyle{\pm 0.051}$ 6 $0.660\scriptstyle{\pm 0.016}$ $0.674\scriptstyle{\pm 0.021}$ $0.667\scriptstyle{\pm 0.008}$ $0.470\scriptstyle{\pm 0.115}$ 7 $0.637\scriptstyle{\pm 0.014}$ $0.677\scriptstyle{\pm 0.009}$ $0.675\scriptstyle{\pm 0.012}$ $0.518\scriptstyle{\pm 0.036}$ 8 $0.667\scriptstyle{\pm 0.016}$ $0.670\scriptstyle{\pm 0.010}$ $0.675\scriptstyle{\pm 0.029}$ $0.446\scriptstyle{\pm 0.030}$ 9 $0.584\scriptstyle{\pm 0.086}$ $0.663\scriptstyle{\pm 0.005}$ $0.610\scriptstyle{\pm 0.104}$ $0.462\scriptstyle{\pm 0.113}$ 10 $0.663\scriptstyle{\pm 0.035}$ $0.674\scriptstyle{\pm 0.029}$ $0.668\scriptstyle{\pm 0.017}$ $0.531\scriptstyle{\pm 0.064}$ 11 $0.648\scriptstyle{\pm 0.017}$ $0.653\scriptstyle{\pm 0.015}$ $0.529\scriptstyle{\pm 0.096}$ $0.484\scriptstyle{\pm 0.047}$ 12 $0.660\scriptstyle{\pm 0.016}$ $0.662\scriptstyle{\pm 0.016}$ $0.684\scriptstyle{\pm 0.019}$ $0.549\scriptstyle{\pm 0.071}$ 13 $0.664\scriptstyle{\pm 0.014}$ $0.655\scriptstyle{\pm 0.017}$ $0.526\scriptstyle{\pm 0.105}$ $0.487\scriptstyle{\pm 0.040}$ 14 $0.650\scriptstyle{\pm 0.011}$ $0.654\scriptstyle{\pm 0.006}$ $0.668\scriptstyle{\pm 0.030}$ $0.505\scriptstyle{\pm 0.039}$ 15 $0.635\scriptstyle{\pm 0.038}$ $0.651\scriptstyle{\pm 0.018}$ $0.673\scriptstyle{\pm 0.009}$ $0.458\scriptstyle{\pm 0.143}$ 16 $0.652\scriptstyle{\pm 0.026}$ $0.654\scriptstyle{\pm 0.007}$ $0.616\scriptstyle{\pm 0.108}$ $0.486\scriptstyle{\pm 0.012}$ 17 $0.628\scriptstyle{\pm 0.019}$ $0.633\scriptstyle{\pm 0.015}$ $0.675\scriptstyle{\pm 0.030}$ $0.518\scriptstyle{\pm 0.057}$ 18 $0.652\scriptstyle{\pm 0.009}$ $0.605\scriptstyle{\pm 0.021}$ $0.598\scriptstyle{\pm 0.109}$ $0.519\scriptstyle{\pm 0.031}$ 19 $0.660\scriptstyle{\pm 0.012}$ $0.581\scriptstyle{\pm 0.037}$ $0.628\scriptstyle{\pm 0.116}$ $0.487\scriptstyle{\pm 0.050}$ 20 $0.651\scriptstyle{\pm 0.013}$ $0.480\scriptstyle{\pm 0.074}$ $0.536\scriptstyle{\pm 0.116}$ $0.560\scriptstyle{\pm 0.043}$ CIFAR-10 CreWra CreEnsα CreBNN CreNet Member $\bm{m}$ 1 $0.942\scriptstyle{\pm 0.001}$ $0.955\scriptstyle{\pm 0.000}$ $0.875\scriptstyle{\pm 0.004}$ $0.941\scriptstyle{\pm 0.000}$ 2 $0.944\scriptstyle{\pm 0.001}$ $0.954\scriptstyle{\pm 0.000}$ $0.877\scriptstyle{\pm 0.005}$ $0.943\scriptstyle{\pm 0.000}$ 3 $0.944\scriptstyle{\pm 0.001}$ $0.952\scriptstyle{\pm 0.001}$ $0.867\scriptstyle{\pm 0.015}$ $0.944\scriptstyle{\pm 0.001}$ 4 $0.943\scriptstyle{\pm 0.002}$ $0.952\scriptstyle{\pm 0.000}$ $0.881\scriptstyle{\pm 0.003}$ $0.942\scriptstyle{\pm 0.001}$ 5 $0.943\scriptstyle{\pm 0.001}$ $0.952\scriptstyle{\pm 0.001}$ $0.872\scriptstyle{\pm 0.004}$ $0.942\scriptstyle{\pm 0.001}$ 6 $0.943\scriptstyle{\pm 0.000}$ $0.952\scriptstyle{\pm 0.000}$ $0.869\scriptstyle{\pm 0.006}$ $0.942\scriptstyle{\pm 0.002}$ 7 $0.941\scriptstyle{\pm 0.001}$ $0.952\scriptstyle{\pm 0.001}$ $0.877\scriptstyle{\pm 0.001}$ $0.943\scriptstyle{\pm 0.001}$ 8 $0.942\scriptstyle{\pm 0.001}$ $0.951\scriptstyle{\pm 0.001}$ $0.880\scriptstyle{\pm 0.006}$ $0.943\scriptstyle{\pm 0.002}$ 9 $0.943\scriptstyle{\pm 0.001}$ $0.951\scriptstyle{\pm 0.000}$ $0.879\scriptstyle{\pm 0.005}$ $0.944\scriptstyle{\pm 0.001}$ 10 $0.943\scriptstyle{\pm 0.000}$ $0.953\scriptstyle{\pm 0.000}$ $0.873\scriptstyle{\pm 0.007}$ $0.943\scriptstyle{\pm 0.003}$ 11 $0.942\scriptstyle{\pm 0.001}$ $0.953\scriptstyle{\pm 0.001}$ $0.882\scriptstyle{\pm 0.004}$ $0.943\scriptstyle{\pm 0.000}$ 12 $0.943\scriptstyle{\pm 0.000}$ $0.953\scriptstyle{\pm 0.000}$ $0.879\scriptstyle{\pm 0.003}$ $0.942\scriptstyle{\pm 0.000}$ 13 $0.943\scriptstyle{\pm 0.000}$ $0.953\scriptstyle{\pm 0.001}$ $0.874\scriptstyle{\pm 0.000}$ $0.943\scriptstyle{\pm 0.002}$ 14 $0.945\scriptstyle{\pm 0.000}$ $0.952\scriptstyle{\pm 0.001}$ $0.872\scriptstyle{\pm 0.006}$ $0.941\scriptstyle{\pm 0.001}$ 15 $0.942\scriptstyle{\pm 0.002}$ $0.948\scriptstyle{\pm 0.000}$ $0.856\scriptstyle{\pm 0.019}$ $0.942\scriptstyle{\pm 0.001}$ 16 $0.942\scriptstyle{\pm 0.000}$ $0.941\scriptstyle{\pm 0.000}$ $0.870\scriptstyle{\pm 0.002}$ $0.942\scriptstyle{\pm 0.001}$ 17 $0.942\scriptstyle{\pm 0.001}$ $0.930\scriptstyle{\pm 0.001}$ $0.854\scriptstyle{\pm 0.035}$ $0.942\scriptstyle{\pm 0.002}$ 18 $0.944\scriptstyle{\pm 0.001}$ $0.921\scriptstyle{\pm 0.001}$ $0.872\scriptstyle{\pm 0.007}$ $0.943\scriptstyle{\pm 0.002}$ 19 $0.943\scriptstyle{\pm 0.001}$ $0.906\scriptstyle{\pm 0.002}$ $0.873\scriptstyle{\pm 0.001}$ $0.942\scriptstyle{\pm 0.001}$ 20 $0.942\scriptstyle{\pm 0.001}$ $0.872\scriptstyle{\pm 0.002}$ $0.860\scriptstyle{\pm 0.012}$ $0.942\scriptstyle{\pm 0.001}$ QualityMRI CreWra CreEnsα CreBNN CreNet Member $\bm{m}$ 1 $0.457\scriptstyle{\pm 0.099}$ $0.430\scriptstyle{\pm 0.170}$ $0.548\scriptstyle{\pm 0.115}$ $0.581\scriptstyle{\pm 0.137}$ 2 $0.452\scriptstyle{\pm 0.139}$ $0.409\scriptstyle{\pm 0.163}$ $0.484\scriptstyle{\pm 0.115}$ $0.559\scriptstyle{\pm 0.112}$ 3 $0.425\scriptstyle{\pm 0.124}$ $0.419\scriptstyle{\pm 0.160}$ $0.468\scriptstyle{\pm 0.126}$ $0.570\scriptstyle{\pm 0.110}$ 4 $0.419\scriptstyle{\pm 0.126}$ $0.446\scriptstyle{\pm 0.187}$ $0.548\scriptstyle{\pm 0.115}$ $0.575\scriptstyle{\pm 0.102}$ 5 $0.398\scriptstyle{\pm 0.135}$ $0.387\scriptstyle{\pm 0.139}$ $0.468\scriptstyle{\pm 0.126}$ $0.548\scriptstyle{\pm 0.126}$ 6 $0.403\scriptstyle{\pm 0.117}$ $0.398\scriptstyle{\pm 0.158}$ $0.554\scriptstyle{\pm 0.118}$ $0.554\scriptstyle{\pm 0.084}$ 7 $0.409\scriptstyle{\pm 0.175}$ $0.414\scriptstyle{\pm 0.161}$ $0.446\scriptstyle{\pm 0.107}$ $0.554\scriptstyle{\pm 0.095}$ 8 $0.430\scriptstyle{\pm 0.119}$ $0.441\scriptstyle{\pm 0.182}$ $0.554\scriptstyle{\pm 0.118}$ $0.559\scriptstyle{\pm 0.133}$ 9 $0.392\scriptstyle{\pm 0.137}$ $0.419\scriptstyle{\pm 0.181}$ $0.473\scriptstyle{\pm 0.110}$ $0.586\scriptstyle{\pm 0.141}$ 10 $0.446\scriptstyle{\pm 0.177}$ $0.430\scriptstyle{\pm 0.140}$ $0.462\scriptstyle{\pm 0.130}$ $0.570\scriptstyle{\pm 0.118}$ 11 $0.425\scriptstyle{\pm 0.137}$ $0.414\scriptstyle{\pm 0.153}$ $0.554\scriptstyle{\pm 0.118}$ $0.602\scriptstyle{\pm 0.107}$ 12 $0.398\scriptstyle{\pm 0.122}$ $0.430\scriptstyle{\pm 0.153}$ $0.554\scriptstyle{\pm 0.118}$ $0.565\scriptstyle{\pm 0.103}$ 13 $0.398\scriptstyle{\pm 0.130}$ $0.430\scriptstyle{\pm 0.140}$ $0.446\scriptstyle{\pm 0.107}$ $0.575\scriptstyle{\pm 0.090}$ 14 $0.409\scriptstyle{\pm 0.066}$ $0.430\scriptstyle{\pm 0.142}$ $0.457\scriptstyle{\pm 0.122}$ $0.581\scriptstyle{\pm 0.070}$ 15 $0.409\scriptstyle{\pm 0.146}$ $0.414\scriptstyle{\pm 0.132}$ $0.468\scriptstyle{\pm 0.126}$ $0.624\scriptstyle{\pm 0.133}$ 16 $0.414\scriptstyle{\pm 0.107}$ $0.425\scriptstyle{\pm 0.167}$ $0.554\scriptstyle{\pm 0.118}$ $0.608\scriptstyle{\pm 0.100}$ 17 $9\scriptstyle{\pm 0.139} 0.414±0.1$$4\scriptstyle{\pm 0.119}$ $0.554\scriptstyle{\pm 0.118}$ $0.581\scriptstyle{\pm 0.091}$ 18 $0.403\scriptstyle{\pm 0.168}$ $0.414\scriptstyle{\pm 0.129}$ $0.468\scriptstyle{\pm 0.126}$ $0.581\scriptstyle{\pm 0.117}$ 19 $0.425\scriptstyle{\pm 0.151}$ $0.414\scriptstyle{\pm 0.073}$ $0.554\scriptstyle{\pm 0.118}$ $0.565\scriptstyle{\pm 0.142}$ 20 $0.430\scriptstyle{\pm 0.154}$ $0.446\scriptstyle{\pm 0.068}$ $0.548\scriptstyle{\pm 0.117}$ $0.581\scriptstyle{\pm 0.103}$